Simplify and hence solve the following equations: (a) 2* (i) In(5x +2) = 2.1 (ii)
Answer #1

22 m-n (aman = M-1) com a => m= n) ga 2x 2-² 24 > x-1/2 = 20 => 2x-x= -1/2 > a=-/ x = -2 La 15x+3) = 21 en (5x+2) - 2 ___2L >